Assume that interest is the only finance charge. Use financial calculator to answer the questions. How much interest would be paid on a $10,000 installment loan to be repaid in 36 monthly installments of $332.14? Round the answer to 4 decimal places. What is the APR on this loan? Round the answer to 2 decimal places.
Answer #1

Interest paid on the loan=332.14*36-10000=1957.04

APR on the loan

Using financial calculator

N=36

PMT=-332.14

PV=10000

FV=0

CPT I/Y=1%

Hence, APR=1%*12=12%